The price p and the demand x for a product are related by the price demand equation x + 300p = 6000. Find the elasticity of demand E(p) interpret E(2)


1
Expert's answer
2021-08-13T13:58:04-0400

"x+300p=6000\\x=6000-300p\\Ed=\frac{dQ}{dP}\times\frac{P}{Q}=-300\times\frac{P}{6000-300p}\\=-300\times\frac{2}{6000-300(2)}\\=-0.11,\space inelastic"